# Genesis: Multimodal Driving Scene Generation with Spatio-Temporal and Cross-Modal Consistency

### v1

@@ 1 Introduction
As shown in Fig. 1, existing driving scene generation methods typically focus on generating data in a single modality, usually RGB videos bevgen ; magicdrive ; drivedreamer ; drivedreamer2 ; drivewm ; panacea ; dive or LiDAR point clouds lidargen ; lidardm ; transfusion ; cameralidar ; deepfusion ; bevfusion . While these approaches have significantly advanced the field of driving scene generation, they overlook the synergistic potential of multimodal generation and lack consistency in aligning RGB videos with various sensor data, leading to limitations in real-world applications. Many of these methods rely on one-step layout-to-data pipelines conditioned solely on coarse spatial priors, such as BEV maps wang2024driving ; street ; magicdrive ; tune ; bevcontrol ; drivedreamer ; drivedreamer2 ; drivewm or 3D boxes kitti ; nuscenes ; waymo , which restricts their ability to capture complex scene dynamics and fine-grained semantics. To propel driving scene generation into the multimodal domain, recent work, such as UniScene li2024uniscene introduces occupancy grids for multimodal generation, but its decoupled design and weak semantic supervision limit cross-modal alignment and scene fidelity. Moreover, like most existing multimodal generation approaches, current methods generally rely on limited semantic supervision—typically in the form of coarse labels or generic captioning models—without fully leveraging the fine-grained descriptive capabilities of modern vision-language models (VLMs) internvl ; liu2023visual ; qwen . This lack of structured semantic grounding restricts the fidelity, controllability, and contextual alignment of generated scenes.
@@ 4.6 Ablation Studies
4.7 Downstream Task Utility
To evaluate the semantic controllability and structural fidelity of generated videos, we assess their utility on downstream segmentation and detection tasks. Performance on these tasks serves as a proxy for the alignment of synthesized content with real-world semantics.
Table 5: Video segmentation and detection.
@@ 4.7 Downstream Task Utility
Table 6: 3D object detection.
@@
Baseline bevfusion C 35.56 41.21
Ours C 34.73 (-0.83) 40.77 (-0.44)
Baseline bevfusion L 64.64 69.23
Ours L 60.82 (-3.82) 67.27 (-1.96)
Baseline bevfusion C&L 68.57 71.46
Ours C&L 65.77 (-2.80) 69.92 (-1.54)
We evaluate the utility of generated data on several downstream perception tasks. As shown in Table 6, our method achieves the best mIoU (38.01) and mAP (27.90) in BEVFormer 3D object detection, surpassing prior methods like DiVE dive and Cogen Cogen , demonstrating strong semantic alignment and visual fidelity.Table 6 presents 3D object detection results using BEVFusion bevfusion . While slight performance drops are observed when training on our generated data (e.g., mAP -2.80 for C&L) , the overall detection performance remains competitive. This indicates that our generated data closely approximates real world data in both quality and fidelity, making it sufficiently realistic and reliable to support model training. These results collectively confirm that our framework produces high-quality multimodal data that is not only visually plausible but also practically valuable for downstream autonomous driving applications.
@@ Appendix A Technical Appendices and Supplementary Material
Figure 9: Controllable generation across time-of-day. By altering scene-level conditions, our method produces consistent multi-view videos aligned with the same underlying map and object layout, while adapting appearance to represent daytime and nighttime settings

### v2

@@ 1 Introduction
As shown in Fig. 1, existing driving scene generation methods typically focus on generating data in a single modality, usually RGB videos bevgen ; magicdrive ; drivedreamer ; drivedreamer2 ; drivewm ; panacea ; dive ; ma2024unleashing or LiDAR point clouds lidargen ; lidardm ; transfusion ; cameralidar ; deepfusion ; bevfusion . While these approaches have significantly advanced the field of driving scene generation, they overlook the synergistic potential of multimodal generation and lack consistency in aligning RGB videos with various sensor data, leading to limitations in real-world applications. Many of these methods rely on one-step layout-to-data pipelines conditioned solely on coarse spatial priors, such as BEV maps street ; magicdrive ; tune ; bevcontrol ; drivedreamer ; drivedreamer2 ; drivewm or 3D boxes kitti ; nuscenes ; waymo , which restricts their ability to capture complex scene dynamics and fine-grained semantics. To propel driving scene generation into the multimodal domain, recent work, such as UniScene li2024uniscene introduces occupancy grids for multimodal generation, but its decoupled design and weak semantic supervision limit cross-modal alignment and scene fidelity. Moreover, like most existing multimodal generation approaches, current methods generally rely on limited semantic supervision—typically in the form of coarse labels or generic captioning models—without fully leveraging the fine-grained descriptive capabilities of modern vision-language models (VLMs) internvl ; liu2023visual ; qwen . This lack of structured semantic grounding restricts the fidelity, controllability, and contextual alignment of generated scenes.
@@ 4.6 Ablation Studies
4.7 Domain gap and Downstream task Utility
Our evaluation approach examines generated videos from two perspectives: the degree of reality measured by domain map metrics, and the functional utility demonstrated in downstream training applications such as 3D object detection tasks.
Table 5: Domain gap on bev segmentation.
@@ 4.7 Domain gap and Downstream task Utility
Table 6: Effect of Multimodal Data Generation on 3D Object Detection. Inputs of the methods in the table are both camera and lidar modalities.
@@
Baseline bevfusion 66.87 69.65
Ours(+cam_gen) 67.09 (+0.22) 70.12 (+0.47)
Ours(+lidar_gen) 67.69(+0.82) 70.58 (+0.93)
Ours(+cam&lidar_gen) 67.78(+0.91) 71.13(+1.48)
As shown in Table 6, we generate videos according to conditions from validation set to metric the domain gap following magicdrivedit . Our method achieves the best mIoU (38.01) and mAP (27.90) on bev segmentation, surpassing prior methods like DiVE dive and Cogen Cogen , demonstrating strong semantic alignment and visual fidelity. As shown in Tab.6, we evaluate the effectiveness of our generative data on the BEVFusionbevfusion framework for 3D object detection. Our approach yields consistent improvements across all settings, increasing the mAP from 66.87 to 67.78 and NDS from 69.65 to 71.13. Notably, joint generation of both camera and LiDAR modalities achieves the highest gains (+0.91 mAP / +1.48 NDS), demonstrating the complementary benefits of multimodal generation. These results validate the utility of high-quality synthetic data in enhancing downstream perception tasks, especially in data-scarce or long-tail scenarios.
@@ Appendix A Technical Appendices and Supplementary Material
Figure 9: Controllable generation across time-of-day. By altering scene-level conditions, our method produces consistent multi-view videos aligned with the same underlying map and object layout, while adapting appearance to represent daytime and nighttime settings.